skip to main content NIST Center for Neutron Research NIST Center for Neutron Research National Institute of Standards and Technology
Home Live Data Instruments CHRNS Proposals
Expansion home Overview Views BT-9 cold source Guides Instrumentation User services Time lines Occasional diary Photos FAQs Contacts

An occasional diary

What follows is a brief description of some of the more visible work that is going on during the long shutdown. The emphasis is on the instruments, the guides, and the new cold source. There is much more going on than what's listed below.
Text in black refers to general matters.
Text in blue refers to the neutron source, the existing cold source, and the guides.
Text in magenta refers to instruments in the confinement building, including the new cold source (PeeWee) that will be installed at BT-9.
Text in green refers to instruments in the guide hall(s).
Text in brown refers to user services.
A partial list of acronyms/abbreviations/special names, may be found at the bottom of the page.
For a plan view that shows the numbering of the beam lines, and a 3-D view of the new guide casings, click on the "Views" tab.



Activities since February 19, 2012
(For a description of activities up until February 19 please click here).

04/26/12 The reactor began a normal 38-day cycle at 20 MW.
04/12/12 The reactor was operated today at 20 MW.
04/04/12 Several instruments in G-100 have been used to collect data: DCS, NG3-SANS, SPINS, NSE, and NG7-SANS.
04/04/12 The reactor was operated today at 10 MW.
03/27/12 The reactor was operated today at 5 MW.
03/27/12 Data have been collected with four of the instruments in C-100: BT-1, BT-4, BT-5, and BT-7.
03/27/12 Effective 03/21 the NCNR has beneficial occupancy of the Secondary Cooling Pump Building and the new Compressor building.
03/26/12 The old and new (East and West) Guide Halls are now combined into a single restricted area for radioactive contamination control.
03/15/12 The reactor was operated today at 1 MW.
03/12/12 Both of the cold sources have been filled with liquid hydrogen and are currently operational.
03/09/12 Hydrogen has been loaded into the Unit 2 cold source and PeeWee.
03/09/12 The MACS (incident beam) crystal filter exchanger (CFX) has been installed at BT-9.
03/08/12 Helium containment has been established surrounding all cold source components.
03/08/12 The stacking of concrete pavers in D-100 has been completed in record time. Congratulations to all involved!
03/06/12 Concrete pavers are being stacked on top of the corrugated steel roof shielding support structure in D-100.
03/05/12 The corrugated steel support structure for the roof shielding in D-100 has been installed.
03/05/12 The NG-A guide casing has been aligned and leak-tested.
03/01/12 The MACS (incident beam) crystal filter exchanger (CFX) has been tested.
03/01/12 The C-100 beam shutters are being tested.
02/29/12 The cold source compressors and refrigerator have been run to purify the helium inventory.
02/29/12 The NG-A guide casing has been installed in D-100.
02/28/12 The refrigerator was operated for the first time since April 4, 2011. It cooled the new load lines and both condensers to their expected operating temperature, 15 K.
02/28/12 The NG-B guide casing has been aligned and leak-tested.
02/27/12 The NG-B guide casing has been installed in D-100.
02/24/12 All scattering instruments in C-100 are available for testing.
02/24/12 All scattering instruments in G-100, other than HFBS, DCS, and NG7SANS, are available for testing.
02/23/12 The unit 2 cold source ballast tank has been successfully pressure tested.
02/23/12 The NG-C guide casing has been installed in D-100. It has also been aligned and leak-tested.
02/21/12 The NG-D guide casing has been installed in D-100. It has since been aligned and leak-tested.



Activities up until February 19, 2012
(For a description of activities since February 19 please click here).

Week 46, February 13-19, 2012:
All of the stanchions for the guides have been mounted in D-100.
The guide loading mechanism has been tested.
C-100 electrical work on the south wall of G-100 is complete.

Week 45, February 6-12, 2012:
The reactor was restarted to low power on Wednesday morning, February 8, 2012.
This was a major milestone on the critical path of the outage.
The walls of the tub shield in D-100 have been poured.
The 8 mm position sensitive detectors for the VSANS instrument have been assembled and connected to the front end electronics.
Electrical power has been restored to the NG-2 backscattering instrument and the tank is being pumped down.


Week 44, January 30 - February 5, 2012:
The PeeWee vacuum system is in service.
The cold source instrumentation wiring has been completed.
A long term test of the Thermal Shield Cooling System has been initiated.
Pucks are being glued to the guides in G-100.
The forms for the walls of the tub shield in D-100 are being installed.
Preparations have been made for commissioning of the new secondary cooling system, including the new building and cooling tower cell.


Week 43, January 23-29, 2012:
The MACS CFX (incident beam cryofilter exchanger) has been received from the manufacturer.
Flex vacuum lines for PeeWee are being installed.
The Experimental Chilled Water (ECW) system in C-100 is operational.
The floor component of the tub shield in D-100 has been poured.
Variable frequency drives for the pumps in the SCPB have been installed.
All electrical circuits have been transferred to permanent power from the new substation.
The first few stanchions for the NG-B, C and D guides in G-100 have been installed near the wall between D-100 and G-100.
New windows and masks have been installed either side of the NG-4 filter cryostat.


Week 42, January 16-22, 2012:
The MACS DFM (double focusing monochromator) cask has been installed.
The DFM cask has been fitted with the DFM itself, the VBA (variable beam aperture), and the ICX (incident beam collimator exchanger).
The first stanchion for the NG-D guide in G-100 has been set.
Integral tests of the PeeWee vacuum jackets have been successfully completed.
The cold box and helium transfer lines have been purged of air and filled with high-purity helium.
The refrigerator is ready for operation.
Work on the hydrogen monitors is proceeding.
The PeeWee ballast tank enclosure is being assembled.


Week 41, January 9-15, 2012:
The PeeWee hydrogen system and its surrounding helium containment are complete amd successfully tested.
The experimental cooling water system has been reinstalled in C-100.
The walkway over the guides in C-100 has been reinstalled.
"Mucking out" of the cooling tower basins has been completed.
The instruments in C-100 are being checked out in order to ensure that they are in good working order before reactor startup.
Similarly, the instruments in the guide hall are being checked out in order to ensure that they are in good working order before reactor startup.
The relocation of electrical circuits for the HFBS instrument has begun.
Several key components of the MACS instrument have been moved into C-100.

Week 40, January 2-8, 2012:
The permanent switchgear for the electrical substation has been set in place.
One of the NG-5 sections has been replaced.
A new guide section for NG-1 has been procured.
Relocation of electrical circuits for the SPINS instrument continues in G-100.
The cold source compressor motors, which had been removed for inspection and cleaning, have been received, set in place and realigned.
The controls for the hydrogen detection system have been installed and tested.


Week 39, December 26 - January 1, 2012:
An 8" concrete floor has been poured in D-100.
A concrete block stub wall has been set in the former doorway between G-100 and the D wing.
The NG-7 guide has been cut, and work on the NG-5 guide has begun.
The PeeWee ballast tank enclosure has been prepared for installation.
The BT-7 floor ramp perimeter has been reestablished.
BT-8 side spacer plates have been installed.
The PLC that operates the PeeWee load line control valves is being programmed.
The MACS analyzer/detector assembly has been moved to the high bay (adjacent to C-100).
... and a Happy New Year!

Week 38, December 19-25:
Forms have been installed in D-100, in preparation for pouring a concrete floor slab.
The PeeWee ballast tank has been installed.
The last PeeWee hydrogen line has been welded and is undergoing vacuum tests.
Work has begun on the PeeWee compressor cooling lines.
The BT-9 shutter has been installed.
The BT-7 and BT-8 shield walls have been reinstalled in C-100.
The wall and roof shields that abut the reinforcement plates attached to the South wall of G-100 have been installed.
Modifications to the NG-7 guide have been initiated.
Merry Christmas ...

Week 37, December 12-18:
The linear position sensitive 3He proportional counters for the VSANS instrument have been received.
The control network linking D-100 to the southern two-thirds of the new guide hall has been completed.
An order of seven more shields for the New Guide system has arrived on site.
All of the cold helium transfer lines that run from the refrigerator to the cold neutron source condensers have been installed.
Five floor reinforcing I-beams have been installed in D-100.
Concrete blocks are being installed at the new compressor building.


Week 36, December 5-11:
About half of the new refrigerator lines, for the old cold source and for PeeWee, have been installed.
The MACS load transfer structure is now in its final position and the warm hydrogen line is complete to the end of the structure.
The holes for the new guide casings, through the wall between D-100 and G-100, have been completed.
The roads by the cooling towers have been paved and opened to traffic.
The bridge crane has been erected in the new Compressor Building.
The SCPB is in the final stages of being piped up, with wiring to follow.


Week 35, November 28 - December 4:
Welds on the warm hydrogen line within the MACS load transfer structure have been completed and tested.
The Experimental Cooling Water system for C-100 has been installed at the north wall

The guides in the casings destined for D-100 are being aligned. NG-D has been completed and NG-B is ready for vacuum testing.
Overhead cranes for the SCPB and for the new Compressor Building have arrived.

Week 34, November 21-27:
Steel reinforcing plates have been bolted to the North face of the wall between D-100 and G-100.
New load lines for both cold sources, to connect the refrigerator to the condensers, have arrived.
The MACS shutter base location system has been installed.
The MACS cover shield has been installed and painted.
Happy Thanksgiving!

Week 33, November 14-20:
All of the shields for the guides in C-100 have been installed.
Steel reinforcing plates have been bolted to the South face of the wall between D-100 and G-100.
The secondary cooling water pumps and other equipment have been placed in the SCPB; interior piping and electrical work continues.

Week 32, November 7-13:
All of the guides and guide casings in C-100 have been installed and aligned.
The entire system, NG-A through NG-D, has been tested for leaks.
Polyethylene block shielding has been installed surrounding the guide casings near the biological shield of the reactor.
The instrumentation group is working on the PeeWee control system.

Week 31, October 31 - November 6:
All of the guide casings in C-100 have been installed, and the guides in all but one of the casings, NG-A(2), have been installed and aligned. NG-A(2) needs three more elements installed and aligned. At that point the entire C-100 Guide System will be pumped down and leak checked. Following a successful leak check, shielding installation can be completed.
The Center Drive Bypass, otherwise known as Off Center Drive, otherwise known as the NCNR Bypass Road, is now open to traffic.
The support structure for the additional cooling tower cell has been erected.
Equipment pads have been poured in the SCPB.


Week 30, October 24-30:
Guide monolith #2, spanning vacuum jackets NG-ABCD(1) and NG-BCD(2), and monolith #3, spanning vacuum jackets NG-BCD(2) and NG-CD(3), have been installed. (For an explanation click here, or click the "Views" tab.)
The D-100 casing for NG-A is being loaded with glass.
Tables have been installed over the guide casings in C-100 to carry polyethylene shielding blocks.
Most of the first coat of asphalt for the "Off Center" Drive has been poured.

Week 29, October 17-23:
Fine tuning of the Thermal Shield Cooling System continues.
The PeeWee D2O lines have been installed.
Steel is going up for the SCPB and for the new compressor building.
The support structure for the new cooling tower is being erected.
Concrete pads for the new electrical substation have been poured.
Guide casings are being installed in C-100.
We have received new sample sticks for the 50 mm and 70 mm ILL cryostats.

Week 28, October 10-16:
Full site acceptance tests for the Thermal Shield Cooling System are continuing.
The helium containment around the PeeWee condenser is complete.
Polyethylene bricks have been hand stacked surrounding the guide casings in the wall between C-100 and D-100.
A small hole has been drilled through the wall separating D-100 and G-100 (the guide hall). This will enable location of reinforcing wall plates either side of the wall.


Week 27, October 3-9:
The shutters for NG-B, -C, and -D have been successfully vacuum tested.
The NG-A shutter has been installed and is being vacuum tested.
The PeeWee condenser vacuum jacket is complete.
The insulating vacuum pump skid is complete and ready to move to C-100 to complete controls.
The SCPB mezzanine floor slab has been poured.
Work has begun to reconnect the load panels to the NG-3 30 meter SANS instrument.

Week 26, September 26 - October 2:
The full scale replacement NBSR Thermal Shield Cooling System has been installed and successfully tested.
This was a major milestone on the critical path of the outage.
Work continues on the plumbing of the PeeWee cold source.
The NG-B shutter has been installed and is being tested.

Week 25, September 19-25:
Roughly half of the warm hydrogen lines have been run from the PeeWee condenser to the ballast tank location.
The NG-C/D shutter has been installed. Helium and air lines have been run to the shutter, and the shutter is undergoing vacuum tests.
Shielding has been hand stacked at the shutter.
The common guide casing for NG-A through NG-D, outside the biological shield, has been installed.
The first half of the SCPB floor slab has been poured.

Week 24, September 12-18:
The penetration jackets for the new guides, NG-A through NG-D, have been aligned, sealed, and configured for a building confinement test.
The in-pile sections of guides NG-1 through NG-4 have been radiographed and the results verify their integrity.
The cover flange and plate for the line-of-sight hole have been installed and sealed in preparation for the building confinement test.
The northernmost wall and roof shields for NG-A through NG-D have been installed.
The welds on the thermal shield tank have passed their air pressure test.
The heavy water cooling assemblies for the PeeWee cold source have been installed, the condenser has been mounted to the reactor face, and the liquid hydrogen transfer line is in place.

Week 23, September 5-11:
The PeeWee cold source has been successfully installed in the BT-9 beam hole.
This was a major milestone on the critical path of the outage.
Casings for three of the new guides have been installed through the wall separating C-100 and D-100.
A line-of-sight hole has been cut through the wall separating C-100 and D-100.
The roof and side shields for the new guides have been removed to be filled with wax and steel shot.


Week 22, August 29 - September 4:
The installation of the PeeWee cold source continues.
Software and hardware development for the control room upgrade continues.
SCPB waterproofing and back filling continues.


Week 21, August 22-28:
The PeeWee cold source assembly has been brought into C-100 and prepared for insertion into the BT-9 beam hole.
Work continues dry fitting the roof and side shields for the new guides.
The reactor's RT-4 facility is being taken out.
Assembly of the thermal shield heat exchanger with new plates is in progress.
Footers have been poured for the new cooling tower.


Week 20, August 15-21:
Shields for the new guides in C-100 (NG-A through NG-D) have been filled with wax and steel shot.
The spacer plate for the new shutters has been installed. The face plate has been attached.
The control room upgrade is underway.
Light poles are being set in the new parking area.

Week 19, August 8-14:
Using flexible stainless steel hoses the existing thermal shield cooling lines in C-100 have been connected to the newly installed stainless steel thermal shield headers.
Welding operations associated with the cold source ballast tank in D-200 are mostly completed.
Shields for the new guides in C-100 (NG-A through NG-D) have been dry-fitted prior to filling with wax and steel shot.
The BT-9 keyway has been cleaned and the beam tube is now ready for PeeWee installation.

Week 18, August 1-7:
Vacuum lines and helium return lines for NG-1 through NG-4 have been installed.
The walls of the new Compressor Building have been poured.
Subassemblies for the PeeWee cold source are being fabricated and a platform for PeeWee installation has been constructed.
The new BT-9 shields have been dry-fitted prior to filling with wax and steel shot.
Disassembly of the new guide mockup has commenced.

Week 17, July 25-31:
A second, significantly larger, rectangular hole has been cut between C-100 and D-100 to accommodate the NG-B, C, and D guides.
Structural steel for the SCPB and for the new Compressor Building has been delivered.
Testing of the vacuum jackets for guides NG-1 through NG-7 has been completed, and the roof shields over NG-5 through NG-7 have been replaced. A new vacuum manifold system for NG-1 through NG-4 is being installed.
The mockup guide installation in the high bay is complete and ready for installation in C-100.


Week 16, July 18-24:
The BT-9 beam tube has been cleaned and prepared to accept the new cold source (PeeWee).
New shields for BT-6 have been filled and readied for installation.
A rectangular hole has been cut between C-100 and D-100 for the NG-A guide penetration.
Shields and a platform have been installed on the West side of the biological shield in order to enable work to proceed on the upper ring header of the thermal shield.
Work has started on installation of forms for the walls of the new Compressor Building.
The vacuum jackets for guides NG-1 through NG-7 have been closed up and are undergoing vacuum testing.

Week 15, July 11-17:
The vacuum jackets for the NG-2, NG-3, and NG-4 guides have been put back in place and the guides reinstalled. The common casing for guides NG-5, NG-6 and NG-7 has been put under vacuum.
The walls of the Secondary Cooling Pump Building have been poured.
Some of the relocated electrical panels in the guide hall have been energized.

Week 14, July 4-10:
The common casing for the NG-5 through NG-7 guides has been reinstalled. Final vacuum testing on all upstream vacuum jackets is complete and all are ready for reinstallation.

Week 13, June 27 - July 3:
The first elements of the guides in the CT-E out-of-pile vacuum casing have been replaced and the casing has been successfully leak tested.
The existing cold source hydrogen line has been connected to the ballast tank.
Drilling operations have begun, through the wall from D-100 to C-100.
Preparations continue for installation of the dedicated cold source PeeWee in the BT-9 beam tube.

Week 12, June 20-26:
The NG-5, 6, and 7 monolithic vacuum casing outside the biological shield, situated at the "CT-E" (Cryogenic Tube East), has been removed.
The old in-pile monolith for NG-5, 6, and 7 has been removed and taken to storage. The brand new replacement in-pile monolith has been successfully installed.


Week 11, June 13-19:
The vacuum jackets of NG-2 and NG-4 have been closed and successfully vacuum tested. NG-1 and NG-3 await testing.
Preparations were made for installation of the new in-pile monolith for NG-5 thru NG-7.
Two pilot holes have been drilled from C-100 through to D-100. One is for the NG-A guide; the other one is for NG-B, C, and D.
Guide hall utilities continue to be restored.
Excavation for the new compressor building has commenced.
Final grading of the new Center Drive road continues.


Week 10, June 6-12:
The first elements of guides NG-1 through NG-4 (just outside the biological shield) have been replaced.
The in-pile guide manifold for the new guides (NG-A through NG-D), situated at the "CT-W" (Cryogenic Tube West), has been successfully installed.
The hydrogen lines for the existing cold source, running from C-100 to D-100/D-200, have been successfully leak tested.
The existing cold source ballast tank has been moved to its new home, in D-200.


Week 9, May 30 - June 5:
The bottom sections of four of the original East wing support columns (numbered 3, 4, 5, and 7 counting from the South end of the guide hall), extending roughly 14 feet to the bottom of the supertruss, have been excised and discarded. This will enable installation of the new guides (A through D) and associated instruments. The removal of column 7 affords the possibility of extending the NG-1 guide into the West wing.

Week 8, May 23-29:
A large section of the concrete wall between the guide hall and D-200 has been removed. This will enable relocation of the existing cold source ballast tank to its new home, in D-200.
Work is proceeding to install the hydrogen lines for the existing cold source, running from C-100 to D-100/D-200.


In future diary entries we shall allude to the East and West wings of the guide hall, rather than to the old and new guide halls.

Week 7, May 16-22:
The wall between the old and new guide halls has been removed.
Work has begun to remove structural components that if left in place would potentially obstruct beam lines entering the new guide hall.

The MACS monochromator shielding has all been removed, and many of the roof shields over the existing guides (NG-1 through NG-7) have been removed.

Week 6, May 9-15:
Removal of the wall between the old and new guide halls is nearly complete. The utilities on the west wall of the original guide hall have mostly been removed.
The BT-9 CO2 seal has been removed. This was a major milestone on the critical path of the outage.
Removal of the MACS monochromator shielding is about 75% complete.


Week 5, May 2-8:
Demolition of the wall between the old and new guide halls continued. The lower windows and the stub wall that formed part of the west wall of the original guide hall have been removed.
The MACS analyzer assembly and sample stage have been removed. Removal of the monochromator shielding has begun.
The first shields for MACS-II@BT-9 have arrived and have been dry fitted.


Week 4, April 25 - May 1:
Demolition of the wall between the old and new guide halls has begun.
The ballast tank for the existing cold source has been removed from the confinement building to a temporary storage location in the old guide hall.
Contractors have begun demolition work on the existing thermal shield cooling system.
The BT-9 collimator has been removed and a temporary shielding plug has been inserted.
New shielding has been installed at the engineering diffractometer, BT-8.

Week 3, April 18-24:
Various tasks have been going on behind the scenes. They include the removal of utilities that would otherwise conflict with the new guides and shields, and preparations for removal of the wall between the old and new guide halls.
All of the hardware for the thermal shield cooling system upgrade has been delivered and is currently stored in the new guide hall.

Week 2, April 11-17:
The BT-9 spectrometer has been dismantled and removed to storage.
All of the existing cold source infrastructure, other than the ballast tank, has been removed from the confinement building.

Week 1, April 4-10:
The NG-1 reflectometer has been lifted almost intact and moved to the West wing of the guide hall. Optical support structures at the front end were also removed. The monochromator (inside the guide shielding) and main shutter (attached to the guide shielding) remain.
The Advanced Neutron Diffractometer/Reflectometer (AND/R) has been disassembled and the epoxy floor has been removed. The shutter, monochromator and slit 1 assemblies remain (all within the guide shielding) and will not be moved until later. The electrical supply will be removed by Plant Division.

The complete inventory of hydrogen has been removed from the existing cold source.

Sunday, April 3, 2011:
The source was shut down at 10:30 p.m.




Acronyms/abbreviations/special names:
  • CT-E and CT-W mean Cryogenic Tube East and Cryogenic Tube West respectively. These tubes view the existing cold source. Guides NG-5, 6, and 7 view the cold source through CT-E. Guides NG-A through NG-D view the cold source through CT-W.
  • C-100 is the confinement building, G-100 is the guide hall, and D-100 is a room between C-100 and G-100. See the "Views" tab (below).
  • SCPB is the Secondary Cooling Pump Building.
  • The abbreviations BT and NG refer to beam tubes (that terminate in C-100) and neutron guides (that terminate in G-100) respectively. See the "Views" tab (below).
  • MACS is the Multi-Axis Crystal Spectrometer that will be relocated to the BT-9 beam tube. MACS acronyms include the following:
    • DFM means double focusing monochromator
    • CFX means cryofilter exchanger
    • VBA means variable beam aperture
    • ICX means incident beam collimator exchanger
  • PeeWee is the dedicated cold source that will be installed in the BT-9 beam tube.
Expansion home Overview Views BT-9 cold source Guides Instrumentation User services Time lines Occasional diary Photos FAQs Contacts

NCNR Expansion Initiative main page.
NCNR home page.
 
Last modified 12-April-2012 by website owner: NCNR (attn: Przemek Klosowski)